Given the following two-commodity system where both commodities are perishable and income (Y), is exogenous 

D1 = -2000 + 7Y -200P1+300P2


S1 = -200 + 500P1 - 100P2


D2 = -1000 +4Y + 200P1 - 100P2


S2 = -800 - 100P1 + 300P2

And that for flow equilibrium, D1 = S1 and D2 = S2


 a. Find the reduced form of the system


 b. Hence find the flow equilibrium values of the endogenous variables when the consumers’ income(Y)

is $9.00


 c. Find the change in the flow equilibrium values that result from a unit change in Y.


1
Expert's answer
2021-04-08T07:26:54-0400

(a)D1=S1

"-2000+7Y-200P_{1}+300P_{2}=-200+500P_{1}-100P_{2}"

"7Y=-200=2000=500P_{1}=200p_{1}-100p_{2}-300p_{2}"

"7y=1800+700p_{1}-400p_{2}"

"y=257\frac{1}{7}+100p_{1}-57\frac{1}{7}p_{2}"


D2=S2

"-100+4y+200p_{1}-100p_{2}=-800-100p_{1}+300p_{2}"

"4y=-800+1000-100p_{1}-200p_{1}+300p_{2}+100p_{2}"

"4y=200-300p_{1}+400p_{2}"

"y=50-75p_{1}+100p_{2}"


(b)

"(9=257.142+100p_{1}-57.142p_{2})75"

"(9=50-75p_{1}+100p_{2})100"

"675=19285.714+7500p_{1}-4285.714p_{2}"

"900=5000-7500p_{1}+714.286p_{2}"

"1575=24285.714+5714.286p_{2}"

"p_{2}=-3.97"


"9=50-75p_{1}+100(-3.97)"

"9=50-75p_{1}-397"

"75p_{1}=9+397-50"

"75p_{1}=356"

"p_{1}=-4.75"

(c)

"y=257\frac{1}{7}+100(-4.75)-57\frac{1}{7}(-3.97)=8.75"

"y=50-75(-4.75)+100(-3.97)=9.25"

"\frac{8.75}{9.25}=0.946"

"=1"